# V-STRONG: Visual Self-Supervised Traversability Learning for Off-road Navigation

### 26 Dec 2023 v1

@@ I INTRODUCTION
Autonomous navigation in off-road environments requires an accurate understanding of the terrain, particularly in identifying areas in the scene which can be traversed by the vehicle. However, contrary to on-road scenarios [1, 2, 3, 4], the notion of traversability in unstructured outdoor settings is much more ambiguous, a priori. For instance, vehicle interaction with terrain features such as ground-vegetation, rocks, and debris is strongly correlated to their size, shape, and material appearance. As such, manual assignment of traversability labels to off-road perception data is non-trivial and prone to error [5]. Furthermore, the complexity of terrain characteristics and the sheer variety of environments make comprehensive data collection of all terrain types a daunting task. Although efforts have been made towards generating annotated, off-road datasets [6, 7, 8, 9, 10], these are generally restricted to a particular set of geospatial locations and seasonal conditions.
@@
Figure 1: (Top) The Polaris RZR off-road vehicle used for our data collection. (Bottom-left) The vehicle is equipped with multiple RGB cameras. (Bottom-middle) Segmentation masks from SAM can disambiguate terrain features. (Bottom-right) Self-supervised learning with mask proposals can achieve robust, fine-grained traversability prediction.
@@
In this paper, we posit that leveraging mask segments for self-supervision offers a simple yet pragmatic method for bootstrapping traversability learning: assuming that image pixels corresponding to the same object or terrain patch should have a similar level of traversability, class-agnostic semantic masks can provide strong priors for self-supervised learning. Specifically, we propose a novel method for pixel-level, contrastive self-supervised learning using SAM and mask-based regularization to address the shortcomings of previous self-supervised methods.
@@ II-A Self-supervised Traversability Learning
Given the aforementioned limitations in labeling traversability for off-road terrains, numerous approaches to learning traversability in a self-supervising fashion have been proposed for off-road autonomy [26, 20, 12, 13, 5]. Such efforts have made use of a variety of sensory modalities to provide a self-supervised learning signal. For instance, Castro et al. [13] used IMU $z$-axis measurements as a traversability score, whereas Seo et al. [27] used a combination of proprioceptive sensors and LiDAR to extract vehicle trajectory traces.
II-B Visual Traversability Learning
Self-supervised learning to predict terrain traversability from visual information has become a recent trend for off-road autonomy [26, 17, 16, 5, 12]. A common approach in this category has been to use autoencoder-based anomaly detection to classify visual terrain features that have not been traversed by the system. Inspired by [19], Schmid et al. [16] propose to aggregate the vehicle footprints and project them into image space to crop out traversed regions. Afterward, an autoencoder [28] is trained with the reconstruction loss only on the traversed regions. This makes the model fail to reconstruct not-traversed areas since they are out-of-distribution features. Thus, at test time, they translate the reconstruction error into the traversability score. As remarked by the authors, this approach can be susceptible to illumination changes and may produce visual artifacts due to the nature of the reconstruction loss.
@@ II-B Visual Traversability Learning
Figure 3: Overview of our method. We first incorporate stereo-depth information to filter out occluded trajectory points and then project the trajectory into image space. Then, positive and negative points are sampled based on the trajectory and SAM-predicted mask information. We apply a pre-trained image encoder along with the traversability decoder that outputs traversability features. Afterward, we extract positive and negative features and apply the trajectory-/mask-based contrastive losses to train the decoder. Additionally, we update our traversability prototype vector using a running-average over positive features. This updated prototype vector is used to calculate the similarity at test time, which will be directly translated into traversability costs. Note that a dashed gray arrow denotes gradient stop, and we do not update our encoder during training.
II-C Contrastive Traversability Learning
Contrastive approaches for terrain traversability have been explored in prior work [15, 5, 26, 19]. In Xue et al. [15], prototype vectors are learned from embedded positive and unlabeled terrain patches based on LiDAR features, which also serve to generate pseudo-labels for an additional supervised classification task. Additionally, other approaches propose to train a network to generate discriminative feature embeddings either with acoustic features [26] or proprioceptive sensor readings [19] coupled with weakly supervised labels learning to estimate traversability with relatively less manual labeling. In comparison, our method only utilizes visual input from RGB cameras and does not require other sensor measurements.
Perhaps closest in spirit to our work is Seo et al. [5], where the authors propose to use Positive-Unlabeled (PU) learning [29] with image-level contrastive learning [30]. The authors adopt a normalizing-flow [31] model and apply a PU learning algorithm for binary classification. Additionally, the contrastive loss [30] is applied to augmented images to encourage the model to have good image representations. However, such a contrastive loss may not be sufficient to provide meaningful information in distinguishing traversable and untraverable areas. On the contrary, we utilize contrastive learning to promote a model to separate traversable and untraversable features in the representation space by sampling positive and negative points within and outside the vehicle trajectories.
II-D Vision Foundation Models
@@ III Method
In this section, we first describe trajectory projection and occlusion handling for positive-label generation, followed by trajectory-level and mask-level contrastive loss definitions leveraging SAM mask predictions. We then elaborate on prototype-vector estimation and conversion to a traversability metric. The overview of our method is illustrated in Fig. 3.
@@ III-A Trajectory Projection and Occlusion Handling
We use vehicle trajectories as a self-supervision signal for learning traversability as shown in Fig. 3. Specifically, with given poses of left/right wheels $P_{t:t+T}$ in the global frame from time $t$ to $t+T$, we project the trajectory into the image space by
@@
We then calculate the contour of the projected left/right trajectories and fill in the contour to complete the trajectory region (Fig. 2 (b)). Finally, we randomly sample positive points $\mathbb{P}_{\text{traj}}$ within the completed trajectory and negative points $\mathbb{N}_{\text{traj}}$ outside the trajectory (Fig. 2 (c)).
@@ III-B Contrastive Learning for Traversability
III-C Trajectory Prototype and Traversability
While training the model with contrastive losses, we estimate a trajectory prototype vector $\mathbf{z}\in\mathbb{R}^{D}$, for converting traversability features to cost at inference time. Specifically, we update the prototype vector $\mathbf{z}$ using exponential moving averaging (EMA),
@@ III-C Trajectory Prototype and Traversability
where $\alpha\in[0,1]$ denotes a momentum value. The prototype vector $\mathbf{z}$ and feature $\mathbf{F}_{s}$ are normalized before and after the EMA operation.
@@ IV-C Implementation details
We adopt the encoder of the ViT-H SAM model as our image encoder and freeze it during training. For training, we use a learning rate of 1e-3 and a batch size of 2 and train the model for 1 epoch. We sample 256/1024 positive/negative points for the trajectory contrastive loss, and sample 512/1024 positive/negative points for the mask contrastive loss. Positive/negative samples on the ego vehicle are excluded. We use $\alpha=0.999$ to update the trajectory prototype vector and $T=300$ for trajectory projection. We set the temperature $\tau=0.05$ and the mask contrastive loss weight $\omega_{\text{mask}}=0.05$. These hyper-parameters (i.e., $\tau$ and $\omega_{\text{mask}}$) have been selected based on the averaged L1 error between predicted costs and manual costs obtained from ground-truth segmentation labels. Note that we project the predicted costs into BEV for hyper-parameter selection since the ground truth labels are in BEV.
@@ IV-D Qualitative Results
The strength of our method becomes clearer when it is applied to off-trail scenarios. Fig. 5 illustrates cost prediction results on CA Hills and Mojave Desert sequences, which contain off-trail images. Baselines struggle to find traversable regions correctly or to predict the lethal object as a high cost. Our approach not only distinguishes traversable regions from lethal objects correctly but also assigns different costs for different semantic objects.
@@ IV-E Quantitative Results
Our datasets Since our datasets do not have labels in image space but in BEV, we project the predicted costs into BEV and inpaint the missing costs using the nearest values. Then, we run a model predictive control (MPC) algorithm to obtain the optimal trajectory based on the predicted costmaps (Fig. 6). Specifically, we adopt model predictive path integral (MPPI) [36] as our MPC algorithm and measure the number of collisions with lethal objects over the number of successful MPPI runs (i.e., Collision Rate). To check collision, we use the ground truth labels in BEV and follow the optimal trajectory obtained from running MPPI on predicted costmaps. To achieve a low collision rate, it is important to find all the lethal objects appearing in the image while finding traversable areas correctly. As reported in Table II, our method outperforms other baselines in LT Murray and CA Hills datasets by a large margin. In the case of the Mojave Desert dataset, all three methods show a similar performance achieving nearly 0 collision rates.
@@ IV-F Generalization to New Environments
Ours w/o mask 0.920 0.945 0.892 0.240 0.068 0.855 0.932
Ours w/ mask 0.959 0.975 0.929 0.100 0.075 0.934 0.925
@@
Figure 9: Qualitative results on the RELLIS-3D dataset with and without the mask loss. Incorporating mask information significantly improves discriminability between traversable and untraversable areas.
@@ IV-G Ablation Study
Effectiveness of using masks Using vehicle trajectories alone for self-supervision is insufficient since the trajectory cannot cover the whole traversable region. This ablation study aims to reveal how mask-based self-supervision is effective in addressing such a concern by comparing results with and without mask-based loss. Table IV presents ablation results when we train the model with and without the mask-based loss. As reported, applying the mask-based loss brings significant improvements in overall metrics. Additionally, as shown in Fig. 9, the model without mask information overfits vehicle trajectories and marks left and right side regions as high costs even though they are indeed traversable. On the other hand, once we apply the mask-based loss, the model correctly identifies traversable regions and marks them as low costs, mitigating the overfitting problem.

### 15 Mar 2024 v2

@@ I INTRODUCTION
Autonomous navigation in off-road environments requires an accurate understanding of the terrain, particularly in identifying traversable areas by the vehicle. However, contrary to on-road scenarios [1, 2, 3, 4], the notion of traversability in unstructured outdoor settings is much more ambiguous, a priori. For instance, vehicle interaction with terrain features such as ground-vegetation, rocks, and debris is strongly correlated to their size, shape, and material appearance. As such, manual assignment of traversability labels to off-road perception data is non-trivial and prone to error [5]. Furthermore, the complexity of terrain characteristics and the sheer variety of environments make comprehensive data collection of all terrain types a daunting task. Although efforts have been made towards generating annotated, off-road datasets [6, 7, 8, 9, 10], these are generally restricted to a particular set of geospatial locations and seasonal conditions.
@@
Figure 1: (Top) The Polaris RZR off-road vehicle used for our data collection. (Bottom-left) The vehicle is equipped with multiple RGB-D cameras. (Bottom-middle) Segmentation masks from SAM can disambiguate terrain features. (Bottom-right) Self-supervised learning with mask proposals can achieve robust, fine-grained traversability prediction.
@@
In this paper, we posit that leveraging mask segments for self-supervision offers a simple yet pragmatic method for bootstrapping traversability learning: assuming that image pixels corresponding to the same object or terrain patch should have a similar level of traversability, class-agnostic semantic masks can provide strong priors for self-supervised learning. Specifically, we propose a novel method for pixel-level, contrastive self-supervised learning using projected vehicle trajectories in pixel space along with mask proposals from SAM, overcoming the limitation of solely using trajectory masks as shown in Fig. 2.
@@ II-A Self-supervised Traversability Learning
Given the aforementioned limitations in labeling traversability for off-road terrains, numerous approaches to learning traversability in a self-supervising fashion have been proposed for off-road autonomy [26, 20, 12, 13, 5, 17, 16]. Such efforts have made use of a variety of sensory modalities to provide a self-supervised learning signal. For instance, Castro et al. [13] used IMU $z$-axis measurements as a traversability score, whereas Seo et al. [27] used a combination of proprioceptive sensors and LiDAR to extract vehicle trajectory traces.
Recently, several approaches [26, 17, 16, 5, 12] have proposed to predict traversability from visual information. Schmid et al. [16] proposed to use autoencoder-based anomaly detection to classify visual terrain features that the system has not traversed. Inspired by [19], they aggregate the vehicle footprints and project them into image space to crop out traversed regions. Then, the autoencoder [28] is trained to reconstruct the traversed regions only. This forces the model to fail on not-traversed areas since they are not observed during the training (i.e., out-of-distribution areas). Thus, at test time, they translate the reconstruction error into the traversability score. However, as the authors remarked, this approach can be susceptible to illumination changes and may produce visual artifacts due to the nature of the reconstruction loss.
@@
Figure 3: Overview of our method. We first incorporate stereo-depth information to filter out occluded trajectory points and then project the trajectory into image space. Then, positive and negative points are sampled based on the trajectory and SAM-predicted mask information. We apply a pre-trained image encoder along with the traversability decoder that outputs traversability features. Afterward, we extract positive and negative features and apply the trajectory-/mask-based contrastive losses to train the decoder. Additionally, we update our traversability vector using a running-average over positive features. This updated vector is used to calculate the similarity at test time, which will be directly translated into traversability costs. Note that a dashed gray arrow denotes gradient stop, and we do not update our encoder during training.
II-B Contrastive Traversability Learning
Several approaches have adopted contrastive learning in their traversability learning [15, 5, 26, 19]. In Xue et al. [15], prototype vectors are learned from embedded positive and unlabeled terrain patches based on LiDAR features, which also serve to generate pseudo-labels for an additional supervised classification task. Additionally, other approaches propose to train a network to generate discriminative feature embeddings either with acoustic features [26] or proprioceptive sensor readings [19] coupled with weakly supervised labels learning to estimate traversability with relatively less manual labeling. In comparison, our method only utilizes visual input from RGB cameras and does not require other sensor measurements.
Perhaps closest in spirit to our work is Seo et al. [5], where the authors propose to use Positive-Unlabeled (PU) learning [29] with image-level contrastive learning [30]. The authors adopt a normalizing-flow [31] model and apply a PU learning algorithm for binary classification. Additionally, the contrastive loss [30] is applied to augmented images to encourage the model to have good image representations. However, such a contrastive loss may not be sufficient to provide meaningful information in distinguishing traversable and non-traversable areas. On the contrary, we utilize contrastive learning to promote a model to separate traversable and non-traversable features in the representation space by sampling positive and negative points within and outside the vehicle trajectories.
II-C Vision Foundation Models
@@ III Method
In this section, we first describe trajectory projection and occlusion handling for positive-label generation, followed by trajectory-level and mask-level contrastive loss definitions leveraging SAM mask predictions. We then elaborate on traversability vector estimation and conversion to a traversability metric. The overview of our method is illustrated in Fig. 3.
@@ III-A Trajectory Projection and Occlusion Handling
We use vehicle trajectories as a self-supervision signal for learning traversability, as shown in Fig. 3. Specifically, with given poses of left/right wheels $P_{t:t+T}$ in the global frame from time $t$ to $t+T$, we project the trajectory into the image space by
@@
Note that the depth estimate $D_{t}$ is obtained from a stereo depth sensor, aligned with the RGB camera. We then calculate the contour of the projected left/right trajectories and fill in the contour to complete the trajectory region (Fig. 2 (b)). Finally, we randomly sample positive points $\mathbb{P}_{\text{traj}}$ within the completed trajectory and negative points $\mathbb{N}_{\text{traj}}$ outside the trajectory (Fig. 2 (c)).
@@ III-B Contrastive Learning for Traversability
III-C Computing Traversability Cost
While training the model with contrastive losses, we estimate a traversability vector $\mathbf{z}\in\mathbb{R}^{D}$, for converting traversability features to cost at inference time. Specifically, we update the traversability vector $\mathbf{z}$ using exponential moving averaging (EMA),
@@ III-C Computing Traversability Cost
where $\alpha\in[0,1]$ denotes a momentum value. The vector $\mathbf{z}$ and feature $\mathbf{F}_{s}$ are normalized before and after the EMA operation. Note that we maintain one traversability vector. This is because the contrastive losses will optimize the traversability decoder to map the SAM feature vectors of traversable regions to have the same representation while distinguishing them from negative feature vectors.
@@ IV-C Implementation details
We adopt the encoder of the ViT-H SAM model as our image encoder and freeze it during training. Note that our full process, from image to prediction, runs around 6 Hz on a single NVIDIA A100 GPU. For training, we use a learning rate of 1e-3 and a batch size of 2 and train the model for 1 epoch. We sample 256/1024 positive/negative points for the trajectory contrastive loss, and sample 512/1024 positive/negative points for the mask contrastive loss. Positive/negative samples on the ego vehicle are excluded. We use $\alpha=0.999$ to update the traversability vector and $T=300$ for trajectory projection. We set the temperature $\tau=0.05$ and the mask contrastive loss weight $\omega_{\text{mask}}=0.05$. These hyper-parameters (i.e., $\tau$ and $\omega_{\text{mask}}$) have been selected based on the averaged L1 error between predicted costs and manual costs obtained from ground-truth segmentation labels. Note that we project the predicted costs into BEV using stereo depths for hyper-parameter selection since the ground truth labels are in BEV.
@@ IV-D Qualitative Results
The strength of our method becomes clearer when it is applied to off-trail scenarios. Fig. 5 illustrates cost prediction results on CA Hills and Mojave Desert sequences, which contain off-trail images. Baselines struggle to find traversable regions correctly or to predict the lethal object as a high cost. Our approach not only distinguishes traversable regions from lethal objects correctly but also assigns different costs for different semantic objects. We conjecture this is because each semantic class has a different frequency to be sampled as positive. For example, the grass can be sampled as positive if the vehicle goes over it, but the tree cannot be sampled because the vehicle cannot pass through it.
@@ IV-E Quantitative Results
Our datasets Since our datasets do not have labels in image space but in BEV, we project the predicted costs into BEV and inpaint the missing costs using the nearest values. Then, we run a model predictive control (MPC) algorithm to obtain the optimal trajectory based on the predicted costmaps (Fig. 6), over all the validation sequences. Specifically, we adopt model predictive path integral (MPPI) [36] as our MPC algorithm and measure the number of collisions with lethal objects over the number of successful MPPI runs (i.e., Collision Rate). To check collision, we use the ground truth labels in BEV and follow the optimal trajectory obtained from running MPPI on predicted costmaps. To achieve a low collision rate, it is important to find all the lethal objects appearing in the image while finding traversable areas correctly. As reported in Table II, our method outperforms other baselines in LT Murray and CA Hills datasets by a large margin. In the case of the Mojave Desert dataset, all three methods show a similar performance achieving nearly 0 collision rates.
@@ IV-F Generalization to New Environments
w/o mask 0.920 0.945 0.892 0.240 0.068 0.855 0.932
w/ mask 0.959 0.975 0.929 0.100 0.075 0.934 0.925
Solely w/ mask 0.955 0.972 0.927 0.129 0.063 0.917 0.937
@@
Figure 9: Qualitative results on the RELLIS-3D dataset with and without the mask loss. Incorporating mask information significantly improves discriminability between traversable and non-traversable areas.
@@ IV-G Ablation Study
Effectiveness of using masks Using vehicle trajectories alone for self-supervision is insufficient since the trajectory cannot cover the whole traversable region. This ablation study aims to reveal how mask-based self-supervision is effective in addressing such a concern by comparing results with and without mask-based loss. As reported in Table IV, applying the mask-based loss brings significant improvements in overall metrics. Additionally, as shown in Fig. 9, the model without mask information overfits vehicle trajectories and marks left and right side regions as high costs even though they are indeed traversable. On the other hand, once we apply the mask-based loss, the model correctly identifies traversable regions and marks them as low costs, mitigating the overfitting problem. Solely using mask-based loss also presents improvements over the one that did not use any mask at all, but it is still behind our final version in most of the metrics. We also observe that false mask proposals from SAM adversely affect the training even though the number of them is very few.
@@ V Conclusion
VI Acknowledgements
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.